# Stein Variational Evolution Strategies

### v1

@@ Appendix A Vanilla CMA-ES Algorithm
3:Sort $\mathbf{y}_{i}$ by $f_{i}$ in ascending order, for $i=1,\dots,m$
$\hat{\mathbf{y}}=\sum_{i=1}^{m}w_{i}\mathbf{y}_{i}$ $\text{where }\sum_{i=1}^{\lambda}w_{i}=1,\forall i.~{}w_{i}>0$ (17)
$\mathbf{x}\leftarrow\mathbf{x}+\alpha_{\mathbf{x}}\sigma\hat{\mathbf{y}}$ (18)
@@
$\mathbf{p}_{\sigma}\leftarrow(1-\alpha_{\sigma})\mathbf{p}_{\sigma}+\sqrt{\alpha_{\sigma}(2-\alpha_{\sigma})~{}\lambda_{\text{eff}}}~{}\mathbf{C}^{-\frac{1}{2}}\hat{\mathbf{y}},$ where $\text{where }\lambda_{\text{eff}}=(\Sigma_{i=1}^{\lambda}w_{i}^{2})^{-1}$
@@
$h_{\sigma}=\mathds{1}\left(\tfrac{\lVert\mathbf{p}_{\sigma}\rVert}{\sqrt{1-(1-\alpha_{\sigma})^{2(t+1)}}}<(1.4+\tfrac{2}{p+1})\operatorname*{\mathbb{E}}\lVert{\cal N}(0,{\rm\bf I})\rVert)\right)$
$\bar{w}_{j}=w_{j}\times\left(1\text{ if }w_{j}\geq 0\text{ else }p/\lVert\mathbf{C}^{-\frac{1}{2}}\mathbf{y}_{j}\rVert^{2}\right)$
@@
$\mathbf{p}_{c}\leftarrow(1-\alpha_{c})\mathbf{p}_{c}+h_{\sigma}\sqrt{\alpha_{c}(2-\alpha_{c})\lambda_{\text{eff}}}~{}\hat{\mathbf{y}}_{i}$
$\mathbf{C}\leftarrow(1+c_{1}d(h_{\sigma})-c_{1}-c_{\lambda}\textstyle\sum_{j=1}^{p}w_{j})\mathbf{C}+c_{1}\mathbf{p}_{c}{\mathbf{p}_{c}}^{T}+c_{\lambda}\sum_{i=1}^{p}\bar{w}_{i}\mathbf{y}_{i}\mathbf{y}_{i}^{T}$
@@ Appendix B Experimental Details
This section lists the full experimental details for this paper. The code is partially based on Langosco di Langosco et al. (2021) and Lange (2023). For the experiments including existing ES such as vanilla CMA-ES, we use the implementations in evosax (Lange, 2023). All the experiments were performed on an internal cluster with eight NVIDIA A40 GPUs. The code to reproduce our experiments and plots will be made available upon conference publication.
Each experiment is repeated 10 times using randomly generated seeds. In each plot, we report the mean performance across all 10 runs and 1.96 standard error bars. To compute the maximum mean discrepancy (Gretton et al., 2012, MMD) we use the RBF kernel for which we select the bandwidth based on the median distance between the ground truth samples (Han and Liu, 2018). For all experiments, we use the standard RBF kernel $k(\mathbf{x},\mathbf{y})=\exp(-\lVert\mathbf{x}-\mathbf{y}\rVert^{2}/2h)$ for which we find the bandwidth $h$ via grid search. Following, Salimans et al. (2017) we use a rank transformation for fitness shaping for all OpenAI-ES-based methods. For GF-SVGD, we follow Han and Liu (2018) in using a Gaussian prior ${\cal N}(0,\sigma^{2}{\rm\bf I})$, where $\sigma$ is determined via a hyperparameter search.
@@ B.1 Hyperparameter Tuning
We tune the hyperparameters for each method separately. For all methods, this includes the kernel bandwidth. For SVGD, GF-SVGD and SV-OpenAI-ES, we additionally tune the Adam learning rate, while the initial step-size $\sigma$ for SV-CMA-ES is selected analogously. For GF-SVGD, we follow Han and Liu (2018) in using a Gaussian prior centered at the origin, with isotropic covariance. The scale of the prior covariance is also determined via grid search. The ranges over which we search are listed in the following subsections in the corresponding hyperparameter paragraph of each experimental subsection. For SV-OpenAI-ES, we additionally tune the width $\sigma$ of the proposal distribution, and for SV-CMA-ES we select the elite population size $\lambda$ via grid search. All remaining hyperparameters for the CMA-ES- and OpenAI-ES-based algorithms are chosen to be the defaults from evosax. The ranges over which we search the hyperparameters differs across the problems due to their different characteristics. We list the specific ranges in the following subsections. The full list of hyperparameters can be found in Table 1.
@@
Task SVGD GF-SVGD SV-CMA-ES SV-OpenAI-ES
Hyperparameter $\epsilon$ $h$ $\epsilon$ $h$ $\sigma^{2}$ $\lambda$ $h$ $\sigma^{2}$ $\epsilon$ $h$ $\sigma^{2}$
Gaussian Mixture 0.05 0.223 1.0 0.889 2.72 2 0.889 0.50 0.50 0.001 0.10
Double Banana 1.0 $10^{-4}$ 0.001 0.011 1.116 2 0.011 0.5 0.001 $10^{-4}$ 0.15
Motion Planning 0.01 0.01 0.001 0.67 2.67 2 0.01 0.10 0.05 0.01 0.10
Covtype 0.005 0.334 0.05 1.0 4.45 9 0.667 0.45 0.01 0.334 0.5
Glass 0.01 1.0 0.1 1.0 4.45 6 1.0 0.40 0.01 1.0 0.05
Pendulum 0.05 16.67 0.34 2 3.334 0.47 0.10 30.0 0.05
CartPole 0.10 13.33 0.45 3 30.0 0.894 1.0 30.0 1.0
MountainCar 0.05 23.33 0.56 2 30.0 0.68 1.0 30.0 0.68
Halfcheetah 0.05 6.667 0.01 5 16.67 0.68 0.05 30.0 0.05
Hopper 0.10 16.67 0.01 5 3.334 0.05 0.10 30.0 0.26
Walker 0.10 10.0 0.01 8 10.0 0.79 0.05 30.0 0.16
@@
GF-SVGD: $\epsilon$ is the Adam learning rate, $h$ is the bandwidth, $\sigma$ is the scale of the prior covariance.
SV-CMA-ES: $\lambda$ is the number of elites, $\sigma$ is the init. step-size, $h$ is the bandwidth.
SV-OpenAI-ES: $\epsilon$ is the Adam learning rate, $\sigma$ is the step-size, $h$ is the bandwidth.
           All values are rounded to three decimal places.
B.2 Sampling from Toy Densities
@@
In our setting, we set $N=4$. We note that one of the sampled weights turned out to be close to zero, which is why Fig. 2 depicts three modes only.
@@
We use the double banana density that was introduced by Detommaso et al. (2018):
@@
We choose the same hyperparameters as Wang et al. (2019a), which are $y=\log(30),\sigma_{1}=1.0,\sigma_{2}=0.09$.
@@
The motion planning problem that we sample from was introduced by Barcelos et al. (2024). The goal is to sample from the density over $N$ waypoints that induce a trajectory from a predefined start point $t_{0}$ to a target $t_{T}$. The resulting density defines a cost that interpolates obstacle avoidance and smoothness penalties:
$p(\mathbf{x})\propto\exp\Bigl{(}\sum_{t=1}^{T}p_{\text{collision}}(\mathbf{x}_{t})+\alpha\lVert\mathbf{x}_{t}-\mathbf{x}_{t-1}\rVert_{2}\Bigr{)},$
@@
The distribution $p_{\text{collision}}$ induces a randomized 2D terrain for navigation, with the valleys between modes having lower cost. We sample a total number of 15 obstacles $\boldsymbol{\mu}_{i}\in{\mathbb{R}}^{2}$ from a Halton sequence following the procedure of Barcelos et al. (2024). We choose the hyperparameters $\sigma=0.25$ and choose $T=5$. This results in an optimization problem over ${\mathbb{R}}^{10}$, as $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ are subject to optimization and each waypoint is two-dimensional.
@@
For each task, we sample 256 ground truth samples to compute the MMD. For the Gaussian Mixtures, samples can be drawn by sampling from the closed form Gaussian mixture. For the other two problems, ground truth samples are drawn with the help of standard Metropolis-Hasting MCMC. For this, we sample from one independent chain per sample, using $100{,}000$ burn-in steps.
@@
We tune the hyperparameters of each method by carrying out a grid search over the following hyperparameter ranges. For the Adam learning rate, we tune over the values proposed by Wang et al. (2019a): $[0.001,0.005,0.01,0.05,0.1,0.5,1.0]$. For the kernel parameter $h$, we use 10 equidistant values from the intervals $[0.001,1]$, $[0.0001,0.1]$, $[0.001,3.0]$ for the Gaussian mixture, double banana and motion planning tasks respectively. For the elite ratio hyperparameter in SV-CMA-ES, we use 10 equidistant values from the interval $[0.15,0.5]$, and for the initial value of $\sigma$ that is used in SV-CMA-ES and SV-OpenAI-ES, we use 10 equidistant values from the interval $[0.05,0.5]$. For GF-SVGD $\sigma$ refers to the prior covariance scaling which we search over 10 equidistant values from the intervals $[0.1,6.0]$, $[0.01,2.0]$, and $[0.01,4.0]$ for the Gaussian mixture, double banana and motion planning tasks respectively.
@@ B.3 Logistic Regression
We follow the implementation of Langosco di Langosco et al. (2021), which uses a hierarchical prior $p(\theta)$ on the parameters $\theta=[\alpha,\beta]$, where $\beta\sim{\cal N}(0,\alpha^{-1})$ and $\alpha\sim\Gamma(a_{0},b_{0})$. Following the aforementioned work, we use $a_{0}=1$ and $b_{0}=0.01$. Given this prior, the task is to approximate samples from the posterior $p(\theta\mid D)=p(D\mid\theta)p(\theta)$ with:
@@
Following the procedure described by Han and Liu (2018), we use the binary version of the Glass dataset. [^2] https://archive.ics.uci.edu/dataset/42/glass+identification For the second logistic regression experiment, we use the binary version of the Covtype dataset following Liu and Wang (2016). [^3] https://archive.ics.uci.edu/dataset/31/covertype This dataset groups the samples into the two groups window and non-window for classification. We partition the data into 70% for training, 10% for validation and 20% for testing and use a batch size of 128.

### v2

@@ Acknowledgements.
3:Sort $\mathbf{y}_{i}$ by $f_{i}$ in ascending order, for $i=1,\dots,n$
$\Delta\mathbf{x}_{\textsc{cma}}=\sigma\sum_{i=1}^{m}w_{i}\mathbf{y}_{i}$ $\text{where }\sum_{i=1}^{m}w_{i}=1,\text{ and }\forall i\in\{1,\dots,m\}.\leavevmode\nobreak\ w_{i}>0$
$\mathbf{x}\leftarrow\mathbf{x}+\epsilon\leavevmode\nobreak\ \Delta\mathbf{x}_{\textsc{cma}}$
@@
$\mathbf{p}_{\sigma}\leftarrow(1-\alpha_{\sigma})\mathbf{p}_{\sigma}+\sqrt{\alpha_{\sigma}(2-\alpha_{\sigma})\leavevmode\nobreak\ m_{\text{eff}}}\leavevmode\nobreak\ \mathbf{C}^{-\frac{1}{2}}\Delta\mathbf{x}_{\textsc{cma}}/\sigma,$ where $\text{where }m_{\text{eff}}=(\textstyle\sum_{i=1}^{m}w_{i}^{2})^{-1}$
@@
$h_{\sigma}=\mathds{1}\left(\tfrac{\lVert\mathbf{p}_{\sigma}\rVert}{\sqrt{1-(1-\alpha_{\sigma})^{2(t+1)}}}<(1.4+\tfrac{2}{d+1})\operatorname*{\mathbb{E}}\lVert{\cal N}(0,{\rm\bf I})\rVert)\right)$
$\bar{w}_{j}=w_{j}\times\left(1\text{ if }w_{j}\geq 0\text{ else }d/\lVert\mathbf{C}^{-\frac{1}{2}}\mathbf{y}_{j}\rVert^{2}\right)$
@@
$\mathbf{p}_{c}\leftarrow(1-\alpha_{c})\mathbf{p}_{c}+h_{\sigma}\sqrt{\alpha_{c}(2-\alpha_{c})m_{\text{eff}}}\leavevmode\nobreak\ \Delta\mathbf{x}_{\textsc{cma}}/\sigma$
$\mathbf{C}\leftarrow(1+\alpha_{1}d(h_{\sigma})-\alpha_{1}-\alpha_{m}\textstyle\sum_{i=1}^{n}w_{i})\mathbf{C}+\alpha_{1}\mathbf{p}_{c}{\mathbf{p}_{c}}^{T}+\alpha_{m}\sum_{i=1}^{n}\bar{w}_{i}\mathbf{y}_{i}\mathbf{y}_{i}^{T}$
Appendix A Vanilla CMA-ES Algorithm
The listing in Algorithm 2 displays the update step of vanilla CMA-ES at iteration $t$. Parallel CMA-ES, on which we base our method, uses the same update, but carries it out $\varrho$ times in parallel, once for each search distribution. We would like to point out that our notation differs from the standard notation of the ES community. While the number of sampled proposals in each iteration is classically denoted by $\lambda$ and the number of selected elites by $\mu$ [Hansen, 2016], we deviate from this notation in the present work. Following the main paper, we denote the number of sampled candidates by $n\in{\mathbb{N}}^{+}$, and the number of elites by $m$. In our experiments, we tune the elite population size $m\in{\mathbb{N}}^{+}$, the initial value for $\sigma$ using a grid search (cf. Appendix B) and use the defaults from the evosax codebase [Lange, 2023] for the remaining parameters. These hyperparameters should coincide with those of Hansen [2016].
A.1 Computation of the Recombination Weights and Hyperparameters
For completeness, we list the computation of the recombination weights and all remaining hyperparameters in this section. Each population uses the same recombination weights, so $w_{ik}=w_{jk}$ for populations $i$ and $j$. For simplicity, we therefore drop the population indices in the weights and define them for the single population case. For this, we use the equations (49)-(58) from Hansen [2016] with an elite population size of $m$. Please note that the aforementioned work uses different notation and denotes the population size by $\lambda$, and the number of elites by $\mu$.
Given population size $n$ and problem dimensionality $d$, we compute the set of weights $w_{1},\ldots,w_{n}$ as follows:
$\epsilon=1$ (16)
$w^{\prime}_{i}=\ln\left(\frac{n+1}{2}\right)-\ln(i),\quad\text{for }i=1,\ldots,n$ (17)
[…]
$m_{\text{eff}}^{-}=\frac{(\sum_{i=m+1}^{n}w^{\prime}_{i})^{2}}{\sum_{i=m+1}^{n}{w^{\prime}}_{i}^{2}}$ (19)
$\alpha_{\sigma}=\frac{m_{\text{eff}}+2}{d+m_{\text{eff}}+5}$ (20)
$d_{\sigma}=1+2\max\Bigl{(}0,\sqrt{\frac{m_{\text{eff}}-1}{d+1}}-1\Bigr{)}+\alpha_{\sigma}$ (21)
$\alpha_{c}=\frac{4+m_{\text{eff}}/d}{d+4+2m_{\text{eff}}/d}$ (22)
$\alpha_{1}=\frac{2}{(d+1.3)^{2}+m_{\text{eff}}}$ (23)
$\alpha_{m}=\min\Bigl{(}1-\alpha_{1},2\frac{1/4+m_{\text{eff}}+1/m_{\text{eff}}-2}{(d+2)^{2}+m_{\text{eff}}}\Bigr{)}$ (24)
$\beta_{m}=1+\alpha_{1}/\alpha_{m}$ (25)
$\beta_{m_{\text{eff}}}=1+\frac{2m_{\text{eff}}^{-}}{m_{\text{eff}}+2}$ (26)
$\beta_{\text{pos def}}=\frac{1-\alpha_{1}-\alpha_{m}}{d\alpha_{m}}$ (27)
$w_{i}=\begin{cases}\frac{1}{\sum_{i=1}^{n}\max(w^{\prime}_{i},0)}w^{\prime}_{i}&\text{if }w^{\prime}_{i}\geq 0\\[5.69054pt] \frac{\min(\beta_{m},\beta_{m_{\text{eff}}},\beta_{\text{pos def}})}{\sum_{i=1}^{n}|\min(w^{\prime}_{i},0)|}w^{\prime}_{i}&\text{else}\end{cases}$ (28)
@@ Appendix B Experimental Details
This section lists the full experimental details for this paper. The code is partially based on Langosco et al. [2021] and Lange [2023]. For the experiments including existing ES such as vanilla CMA-ES, we use the implementations in evosax [Lange, 2023]. All the experiments are performed on an internal cluster with eight NVIDIA A40 GPUs. The code to reproduce our experiments and plots will be made available upon conference publication.
Each experiment is repeated 10 times using randomly generated seeds. In each plot, we report the mean performance across all 10 runs and 1.96 standard error bars. To compute the maximum mean discrepancy [Gretton et al., 2012, MMD] we use the RBF kernel for which we select the bandwidth based on the median distance between the ground truth samples [Han and Liu, 2018]. For all experiments, we use the standard RBF kernel $k(\mathbf{x},\mathbf{y})=\exp(-\lVert\mathbf{x}-\mathbf{y}\rVert^{2}/2h)$ for which we find the bandwidth $h$ via grid search. Following Salimans et al. [2017] we use a rank transformation for fitness shaping for all OpenAI-ES-based methods. For GF-SVGD, we follow Han and Liu [2018] in using a Gaussian prior ${\cal N}(0,\sigma^{2}{\rm\bf I})$, where $\sigma^{2}$ is determined via grid search.
@@ B.1 Hyperparameter Tuning
We tune the hyperparameters for each method separately. For all methods, this includes the kernel bandwidth. For SVGD, GF-SVGD and SV-OpenAI-ES, we additionally tune the Adam learning rate, while the initial step-size $\sigma$ for SV-CMA-ES is selected analogously. For GF-SVGD, we follow Han and Liu [2018] in using a Gaussian prior centered at the origin, with isotropic covariance. The scale of the prior covariance is also determined via grid search. The ranges over which we search are listed in the following subsections in the corresponding hyperparameter paragraph of each experimental subsection. For SV-OpenAI-ES, we additionally tune the width $\sigma$ of the proposal distribution, and for SV-CMA-ES we select the elite population size $m$ via grid search. All remaining hyperparameters for the CMA-ES- and OpenAI-ES-based algorithms are chosen to be the defaults from evosax. The ranges over which we search the hyperparameters differs across the problems due to their different characteristics. We list the specific ranges in the following subsections. The full list of hyperparameters can be found in Table 2.
@@
Task Dim. SVGD GF-SVGD SV-CMA-ES SV-OpenAI-ES
Hyperparameter $\epsilon$ $h$ $\epsilon$ $h$ $\sigma^{2}$ $m$ $h$ $\sigma^{2}$ $\epsilon$ $h$ $\sigma^{2}$
Gaussian Mixture 2 0.05 0.223 1.0 0.889 2.72 2 0.889 0.50 0.50 0.001 0.10
Double Banana 2 1.0 $10^{-4}$ 0.001 0.011 1.116 2 0.011 0.5 0.001 $10^{-4}$ 0.15
Motion Planning 10 0.01 0.01 0.001 0.67 2.67 2 0.01 0.10 0.05 0.01 0.10
Credit 22 0.1 $10^{-3}$ 0.01 0.67 3.34 9 $10^{-3}$ 0.35 0.005 $10^{-3}$ 0.05
Covtype 55 0.01 0.45 0.05 1.0 2.28 12 0.78 0.15 0.01 0.334 0.2
Spam 58 0.01 0.11 0.05 1.0 0.56 9 0.45 0.20 0.01 0.11 0.2
Pendulum 353 0.05 16.67 0.34 2 3.33 0.47 0.10 30.0 0.05
CartPole 386 0.10 13.33 0.45 3 30.0 0.89 1.0 30.0 1.0
MountainCar 337 0.05 23.33 0.56 2 30.0 0.68 1.0 30.0 0.68
Halfcheetah 678 0.05 6.67 0.01 5 16.67 0.68 0.05 30.0 0.05
Hopper 515 0.10 16.67 0.01 5 3.33 0.05 0.10 30.0 0.26
Walker 662 0.10 10.0 0.01 8 10.0 0.79 0.05 30.0 0.16
@@
GF-SVGD: $\epsilon$ is the Adam learning rate, $h$ is the bandwidth, $\sigma^{2}$ is the scale of the prior covariance.
SV-CMA-ES: $m$ is the number of elites, $\sigma^{2}$ is the init. step-size, $h$ is the bandwidth.
SV-OpenAI-ES: $\epsilon$ is the Adam learning rate, $\sigma^{2}$ is the step-size, $h$ is the bandwidth.
B.2 Sampling from Synthetic Densities
@@
In our setting, we set $N=4$. We note that in our experiments one of the sampled weights is close to zero, which is why Fig. 2 depicts only three modes.
@@
We use the double banana density that was introduced by Detommaso et al. [2018]:
@@
We choose the same hyperparameters as Wang et al. [2019a], which are $y=\log(30),\sigma_{1}=1.0,\sigma_{2}=0.09$.
@@
The motion planning problem that we sample from was introduced by Barcelos et al. [2024]. The goal is to sample from the density over $N$ waypoints that induce a trajectory from a predefined start point $t_{0}$ to a target $t_{T}$. The resulting density defines a cost that interpolates obstacle avoidance and smoothness penalties:
$p(\mathbf{x})\propto\exp\Bigl{(}\sum_{t=1}^{W}p_{\text{collision}}(\mathbf{x}_{t})+\alpha\lVert\mathbf{x}_{t}-\mathbf{x}_{t-1}\rVert_{2}\Bigr{)},$
@@
The distribution $p_{\text{collision}}$ induces a randomized 2D terrain for navigation, with the valleys between modes having lower cost. We sample a total number of 15 obstacles $\boldsymbol{\mu}_{i}\in{\mathbb{R}}^{2}$ from a Halton sequence following the procedure of Barcelos et al. [2024]. We choose the hyperparameters $\sigma=0.25$ and choose $N=5$. This results in an optimization problem over ${\mathbb{R}}^{10}$, as each waypoint is two-dimensional.
@@
For each task, we sample 256 ground truth samples to compute the MMD. For the Gaussian Mixtures, samples can be drawn by sampling from the closed form Gaussian mixture. For the other two problems, ground truth samples are drawn with the help of standard Metropolis-Hastings MCMC. For this, we sample from one independent chain per sample, using $100{,}000$ burn-in steps.
@@
We carefully tune the hyperparameters of each method by carrying out a grid search over the following hyperparameter ranges. For the Adam learning rate, we tune over the values proposed by Wang et al. [2019a]: $[0.001,0.005,0.01,0.05,0.1,0.5,1.0]$. For the kernel parameter $h$, we use 10 equidistant values from the intervals $[0.001,1]$, $[0.0001,0.1]$, $[0.001,3.0]$ for the Gaussian mixture, double banana and motion planning tasks respectively. For the elite ratio hyperparameter in SV-CMA-ES, we use 10 equidistant values from the interval $[0.15,0.5]$, and for the initial value of $\sigma$ that is used in SV-CMA-ES and SV-OpenAI-ES, we use 10 equidistant values from the interval $[0.05,0.5]$. For GF-SVGD $\sigma$ refers to the prior covariance scaling which we search over 10 equidistant values from the intervals $[0.1,6.0]$, $[0.01,2.0]$, and $[0.01,4.0]$ for the Gaussian mixture, double banana and motion planning tasks respectively.
@@ B.3 Logistic Regression
We follow the implementation of Langosco et al. [2021], which uses a hierarchical prior $p(\theta)$ on the parameters $\theta=[\alpha,\beta]$, where $\beta\sim{\cal N}(0,\alpha^{-1})$ and $\alpha\sim\Gamma(a_{0},b_{0})$. Following the aforementioned work, we use $a_{0}=1$ and $b_{0}=0.01$. Given data $D$, the task is to approximate samples from the posterior
$p(\theta\mid D)=p(D\mid\theta)p(\theta)$
with
@@
We use three datasets from the UCI Machine Learning repository [Kelly et al., 2023] in our experiments: Covtype [Blackard, 1998], Spambase [Hopkins et al., 1999], and Credit score data [LMU, 2010]. For the Covtype dataset, we create a binary version following Liu and Wang [2016], which groups the samples into the two groups: Lodgepole Pine versus rest. For each task, we partition the data into 70% for training, 10% for validation and 20% for testing and use a batch size of 128. The dimensionalities of the problems are defined by the number of features in the respective datasets, with an additional degree of freedom for the parameter $\alpha$ of the prior.